An aircraft company uses rivets at an approximate demand rate of 2,500 kg per year. Each unit costs $ 30 per kg and the company personnel estimate that it costs $130 to place an order and that the carrying cost of inventory is 10 % per year. How frequently should orders for rivets be placed? Determine the optimum size of each order?

Find EOQ, Order frequency T, Number of orders N and TC
Answer #1

Data: λ(D) = 2,500 Kg. per year

C3 = Rs. 130/-

i = 10%

H = i * c = 30 * 10% = 3

p = Rs. 30/- per unit. (C1 = i × p = 0.10 × 30 = Rs. 3/-)

qo = / ip

qo = / 0.10 × 30 = App. 466 units.

to = qo = / = 466 / 2500 = 0.18 year = 0.18 × 12 = 2.16 month.

N = Number of orders = / qo = 2500 / 466 = 5 orders per year.

EOQ = sqrt(2DS/H)

= sqrt(2*2500*130/3)

= sqrt(216666.67)

= 465.47 units

Order Frequency = R / EOQ

= 2500 / 465.47

= 5.37 orders

Numer Orders is same as order frequency.